OpenStudy (anonymous):

log(x^3 − 125) − log(x^2 + 5x + 25)

OpenStudy (mathstudent55):

\(\log a - \log b = \log \dfrac{a}{b} \) Can you factor those polynomials?

OpenStudy (mathstudent55):

Use the rule of logs above to rewrite this as the log of a fraction. Then factor the numerator and denominator and see if any terms cancel out.

OpenStudy (anonymous):

how do i factor the top one? x^3-125?

Nnesha (nnesha):

\[\huge\rm (x^3 - b^3) = (x-y)(x^2 +xy+y^2)\]

Nnesha (nnesha):